OpenStudy (anonymous):

Hello!!! What comes next in the following sequence? 1, 2, 2, 5, 15, 20, ___ options are: 25,50,100,150. please help with complete explanation

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (moongazer):

25

OpenStudy (anonymous):

Moon, how did you get that answer? Because I don't think it's correct.

OpenStudy (moongazer):

it is a guess ^_^

OpenStudy (anonymous):

Sadaf Manousi, hello.

OpenStudy (anonymous):

Thanks Moon....but the answer at the back of my book is 100...which I can't figure out how?

OpenStudy (anonymous):

Yeah, I think I know how.

OpenStudy (anonymous):

Hi frogwingspan

OpenStudy (anonymous):

Oh wow did I get your name right?

OpenStudy (anonymous):

only first part.

OpenStudy (anonymous):

:<

OpenStudy (anonymous):

Oh wow did I get your name right?

OpenStudy (anonymous):

Yeah, I think I know how.

OpenStudy (anonymous):

Sadaf Manousi, hello.

OpenStudy (anonymous):

Moon, how did you get that answer? Because I don't think it's correct.

OpenStudy (anonymous):

1, 2, 2, 5, 15, 20, ___ 1 = 1 2 = 2 1+2 = 3 1 * 2 = 2 2+3 = 5 5 * 3 = 15 15 + 5 = 20 20 * 5 = 100

OpenStudy (anonymous):

Aw....Thanks alot Suzi :) I think its a logical solution. thanks buddy.

OpenStudy (anonymous):

medal or become fans maybe :D

OpenStudy (anonymous):

both :) just tell me how to do it because I'm new here.

OpenStudy (anonymous):

It's just looking at it really. There is no steadfast technique.

OpenStudy (anonymous):

Oh, you just click her, and click *fan*